# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

### v1

@@ Acknowledgements.
- Siala et al. (2004) Haytham Siala, Robert M O’Keefe, and Kate S Hone. 2004. The impact of religious affiliation on trust in the context of electronic commerce. Interacting with computers 16, 1 (2004), 7–27.
- Spann and Stewart (2018) Alice Spann and Ellen Stewart. 2018. Barriers and facilitators of older people’s mHealth usage: a qualitative review of older people’s views. Human Technology: An Interdisciplinary Journal on Humans in ICT Environments 14, 3 (2018), 264–296.
@@
- Stawarz et al. (2014) Katarzyna Stawarz, Anna L Cox, and Ann Blandford. 2014. Don’t forget your pill! Designing effective medication reminder apps that support users’ daily routines. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 2269–2278.
- Stephan et al. (2017) Laura Siga Stephan, Eduardo Dytz Almeida, Raphael Boesche Guimaraes, Antonio Gaudie Ley, Rodrigo Gonçalves Mathias, Maria Valéria Assis, and Tiago Luiz Luz Leiria. 2017. Processes and recommendations for creating mHealth apps for low-income populations. JMIR mHealth and uHealth 5, 4 (2017), e6510.
@@
- Vajawat et al. (2021) Bhavika Vajawat, Prateek Varshney, and Debanjan Banerjee. 2021. Digital gaming interventions in psychiatry: evidence, applications and challenges. Psychiatry research 295 (2021), 113585.
- Varshney (2014) Upkar Varshney. 2014. Mobile health: Four emerging themes of research. Decision Support Systems 66 (2014), 20–35.
@@
- Vogt and Meier (2010) Joël Vogt and Andreas Meier. 2010. An adaptive user interface framework for eHealth services based on UIML. (2010).
@@
Appendix A Interview and focus group post-session questions
Introduction 2023 Interview/focus group post-session survey Thanks for attending today’s interview/focus group session. We invite you to complete a post-session survey to further share your thoughts on the adaptive user interface for chronic disease-related applications.
Please enter your withdrawal code below (You can
To what extent do you want to have the following main types of adaptation in your daily use of mHealth applications? (Much weaker, Moderately weaker,Slightly weaker, Neutral, Slightly stronger, Moderately stronger, Much stronger)
Presentation adaptation
Content adaptation
Behaviour adaptation
To what extent do you want to have the following presentation adaptation in your daily use of mHealth applications? (Much weaker, Moderately weaker,Slightly weaker, Neutral, Slightly stronger, Moderately stronger, Much stronger)
Graphic design
Color and theme
Font size
Avatar/animation
Information architecture
To what extent do you want to have the following content adaptations in your daily use of mHealth applications (Much weaker, Moderately weaker,Slightly weaker, Neutral, Slightly stronger, Moderately stronger, Much stronger)?
Interface elements rearrangement
Reorder the elements
Hide the elements
Content complexity
Easy to understand language
Minmalize the design
Change text to image
To what extent do you want to have the following behaviour adaptations in your daily use of mHealth applications (Much weaker, Moderately weaker,Slightly weaker, Neutral, Slightly stronger, Moderately stronger, Much stronger)?
Navigation adaptation
Add on functions
Different persuasive strategy
Difficulty levels
Multimodal interaction
Which kind of data do you prefer to use for your interface adaptation?
Your role
Your feedback
Your motivation
Your interaction with the interface
Your game performance
Your physical characteristics
Your psychological characteristics
Your demographic information
Your preference
Your emotions
Your social activity
Your physiological characteristics
your goals
How do you wish your data to be collected?
Analyse of user behaviour through the application
Analysis of activities with keyboard
Smartphone sensor
User input through the application
External sensor
Wearable sensor
What level of initiative do you want to take during the adaptation you prefer?
Manual system
Automatic system
Semi-automatic system
Appendix B Survey questions
Introduction AUIs are software applications where some aspect(s) of the interface is modified to cater for different user needs and/or preferences e.g. font or button size, colour, layout, complexity, interaction style and so on. We are researching how AUIs can be leveraged to better cater for users with chronic diseases. As part of this survey, you will be asked to answer questions about your perspective toward AUI. On average it will take 10-15 minutes for a participant to complete the survey. You can enter a draw for a AU$20 voucher if you complete an online survey. Your email address will not be associated with your response since the contact information is collected in another survey. At the end of the survey, you will also be asked to indicate whether you are interested in participating in a focus group study. You may register through this link if you are interested in the focus group study. If you have any questions, please email wei.wang5@monash.edu. Ethical approval has been provided by Monash University. Further details can be provided upon request.
Survey Consent Form
I have read and understood the Explanatory Statement and I hereby consent to answer the online survey.
I understand that all information I provide for this study will be treated confidentially.
I understand that I am free to contact any of the people involved in the research to seek further clarification and information.
I agree that the data I provide in this research may be used by the investigators in future research projects.
@@ Appendix B Survey questions
Male
Female
Non-binary / third gender
Prefer not to say
Other_
@@
Some high school credit, no diploma, or the equivalent
High school graduate, diploma, or the equivalent
Some college credit, no degree
Trade/technical/vocational training
Associate degree
Bachelor’s degree
Master’s degree
Doctorate degree
Other_
@@
Have you used mHealth apps before to manage your chronic disease? (The most common application of mHealth is the use of mobile devices to educate consumers about preventive healthcare services. However, mHealth is also used for disease surveillance, treatment support, epidemic outbreak tracking and chronic disease management.)
Yes, I have used/curently use the mHealth application
@@
Assistive application help users complete a task or activity more easily
Communication application to communicate with carers, family members, doctors, or emergency contacts
Informative application
Health game application (games that encourage exercise, good living habits)
Health promoting and self-monitoring application
Rehabilitation application
Other (Please specify the name of the application)
@@
Health promoting and self-monitoring application
Rehabilitation application
Other (Please specify the name of the application)
To track what I eat
Help with weight loss
Manage my medications
Increase my physical activity levels
To monitor my chronic disease symptoms
To get education about my chronic disease
Other (Please specify)
How frequently do you use health apps?
less than once a month
Once a week
Less than once a week
2 or more times per day
@@
1-10 minutes
10-20 minutes
more than 30 minutes
more than 60 minutes
@@
Figure 13. Exercise application with AUI
Figure 14. PD-Helper for Parkinson’s Disease Patients Exercise application (Figure 13): The left side is a standard version of the interface, to maintain usability and readability of the interface when the user is running, the interface automatically hide images and changes small-font and enlarge the subtitle and two buttons. (park2015model).
PD-Helper for Parkinson’s Disease Patients (Figure 14): Users are allowed to change the font size and other settings in accordance with their own preferences and capabilities (Jabeen et al., 2019).-
Which kind of adaptation do you prefer for interface adaptation? (mHealth apps can be adapted based on your needs in different ways. Select all that apply)
@@
Which kind of data do you prefer to use for your interface adaptation? (Imagine your mHealth app can be adapted based on your characteristics and preference. Please see different types of data that can be used in adapting mHealth apps. Select all that apply)
@@
Linkedin post
Facebook post/group
Twitter post
Email
Word of mouth
Baker Institute Diabetes Clinic
Alfred Hospital
Dementia Australia, Stroke Foundation, or Kidney Australia
Other social media_

### v2

@@ Acknowledgements.
- Shamsujjoha et al. (2024) Md Shamsujjoha, John Grundy, Hourieh Khalajzadeh, Qinghua Lu, and Li Li. 2024. Developer and End-User Perspectives on Addressing Human Aspects in Mobile eHealth Apps. Information and Software Technology 166 (2024), 107353.
@@
- Sockolow et al. (2021) Paulina S Sockolow, Harleah G Buck, and Efrat Shadmi. 2021. An integrative review of chronic illness mHealth self-care interventions: mapping technology features to patient outcomes. Health informatics journal 27, 3 (2021), 14604582211043914.
@@
- Stowell et al. (2018) Elizabeth Stowell, Mercedes C Lyson, Herman Saksono, Reneé C Wurth, Holly Jimison, Misha Pavel, and Andrea G Parker. 2018. Designing and evaluating mHealth interventions for vulnerable populations: A systematic review. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. 1–17.
- Stoyanov et al. (2015) Stoyan R Stoyanov, Leanne Hides, David J Kavanagh, Oksana Zelenko, Dian Tjondronegoro, and Madhavan Mani. 2015. Mobile app rating scale: a new tool for assessing the quality of health mobile apps. JMIR mHealth and uHealth 3, 1 (2015), e3422.
@@
- Teevan et al. (2009) Jaime Teevan, Susan T Dumais, Daniel J Liebling, and Richard L Hughes. 2009. Changing how people view changes on the web. In Proceedings of the 22nd annual ACM symposium on User interface software and technology. 237–246.
@@
- Todi et al. (2021) Kashyap Todi, Gilles Bailly, Luis Leiva, and Antti Oulasvirta. 2021. Adapting user interfaces with model-based reinforcement learning. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–13.
@@
- Vogt and Meier (2010) Joël Vogt and Andreas Meier. 2010. An adaptive user interface framework for eHealth services based on UIML. In BLED 2010 Proceedings.
- W3C (2015) W3C. 2015. Mobile Accessibility. https://www.w3.org/TR/mobile-accessibility-mapping/
@@
- Yang et al. (2020) Yang Yang, Helen Chen, Hammad Qazi, and Plinio P Morita. 2020. Intervention and evaluation of mobile health technologies in management of patients undergoing chronic dialysis: scoping review. JMIR mHealth and uHealth 8, 4 (2020), e15549.
- Yu et al. (2017) Daihua X Yu, Bambang Parmanto, Brad E Dicianno, Valerie J Watzlaf, and Katherine D Seelman. 2017. Accessibility needs and challenges of a mHealth system for patients with dexterity impairments. Disability and Rehabilitation: Assistive Technology 12, 1 (2017), 56–64.
@@
- Zhou et al. (2020) Leming Zhou, Andi Saptono, I Made Agus Setiawan, and Bambang Parmanto. 2020. Making self-management mobile health apps accessible to people with disabilities: Qualitative single-subject study. JMIR mHealth and uHealth 8, 1 (2020), e15060.
- Zimmerman et al. (2020) John Zimmerman, Changhoon Oh, Nur Yildirim, Alex Kass, Teresa Tung, and Jodi Forlizzi. 2020. UX designers pushing AI in the enterprise: a case for adaptive UIs. Interactions 28, 1 (2020), 72–77.
Appendix A Stage one: User survey questions
Introduction: AUIs are software applications where some aspect(s) of the interface is modified to cater for different user needs and/or preferences, e.g. font or button size, color, layout, complexity, interaction style, and so on. We are researching how AUIs can be leveraged to better cater for users with chronic diseases. As part of this survey, you will be asked to answer questions about your perspective toward AUI. On average it will take 10-15 minutes for a participant to complete the survey. You can enter a draw for a AU$20 voucher if you complete an online survey. Your email address will not be associated with your response since the contact information is collected in another survey. At the end of the survey, you will also be asked to indicate whether you are interested in participating in a focus group study. You may register through this link if you are interested in the focus group study. If you have any questions, please email wei.wang5@monash.edu. Ethical approval has been provided by Monash University. Further details can be provided upon request.
@@ Appendix A Stage one: User survey questions
Have you used mHealth applications before to manage your chronic disease? (The most common application of mHealth is the use of mobile devices to educate consumers about preventive healthcare services. However, mHealth is also used for disease surveillance, treatment support, epidemic outbreak tracking and chronic disease management.)
Yes, I have used/currently use the mHealth application
@@
How frequently do you use health applications?
@@
(a) Exercise application with AUI
(b) PD-Helper for Parkinson’s Disease Patients Figure 16. Two examples of applications with Adaptive User Interfaces (AUI). Exercise application (Figure 16(a)): The left side is a standard version of the interface, to maintain usability and readability of the interface when the user is running, the interface automatically hide images and changes small-font and enlarge the subtitle and two buttons (Park and Lee, 2015).
PD-Helper for Parkinson’s Disease Patients (Figure 16(b)): Users are allowed to change the font size and other settings in accordance with their own preferences and capabilities (Jabeen et al., 2019).-
Which kind of adaptation do you prefer for interface adaptation? (mHealth applications can be adapted based on your needs in different ways. Select all that apply)
@@
Which kind of data do you prefer to use for your interface adaptation? (Imagine your mHealth application can be adapted based on your characteristics and preference. Please see different types of data that can be used in adapting mHealth applications. Select all that apply)
@@
Appendix B Stage two: Guideline evaluation survey
Thank you for taking the time to participate in our survey. This study aims to validate guidelines for designing adaptive user interfaces in mHealth applications by gathering feedback from software practitioners with experience in developing health-related applications. The survey will assess the applicability, clarity, and practicality of the guidelines, helping us determine their potential for integration into real-world development workflows. Important: For the best experience, we recommend completing the survey on a desktop device. If you have any questions, please email wei.wang5@monash.edu. Ethical approval has been provided by Monash University. Further details can be provided upon request.
Section 1: This section is intended to gather basic information about you.
In what age group are you?
To which gender identity do you most identify as?
In which country do you currently reside?
How many employees work in your organisation?
What is your role in the team?(tick all that apply)
Project manager, Business consultant/Marketing manager/Sales personnel, Requirements analyst, Software architect, Programmer, User interface or Graphical User Interface designer/developer/engineer, App animator or operations developer/engineer, QA engineer, Tester, Other (Please specify)
How many years of experience do you have in development, design, or related fields?
Do you have experience developing health-related applications, particularly those focused on chronic diseases?
Yes (Please share details about your experience below.)
Other (Please share details about your experience below.)
Section 2: This section is intended to give you some background information of Adaptive User Interfaces. Details in the User survey • ‣ A
Section 3: Feedback on proposed guidelines. In this section, we aim to gather your thoughts on the guidelines we’ve developed to improve the design of Adaptive User Interfaces in mHealth applications targeting chronic disease. Here is a link to our proposed guidelines. The subsequent questions will ask for your feedback, understanding, and opinions regarding the proposed guidelines. Please review the guidelines thoroughly before answering the following questions and ensure the guidelines are open in another window for reference. We have included brief comprehension check questions in the survey to ensure that all participants are familiar with the guidelines provided.
Based on the guidelines you have reviewed, could you identify which guideline aligns with the following example design? ”Customizations are applied as users log into the application for the first time, modifying the application entirely to suit their specific needs.” If you’re unsure of the answer, please take a moment to reread the guidelines before responding.
Based on the guidelines you have reviewed, could you identify which guideline aligns with the following purpose? ”Enhance the impact of implemented adaptations and engage users seamlessly without interrupting their normal application usage.” If you’re unsure of the answer, please take a moment to reread the guidelines before responding.
Are you able to clearly understand and distinguish between the different guidelines?
How do you perceive the usefulness of each individual guideline?
Would you prefer using our proposed guidelines over existing, standalone guidelines in mHealth applications development process?
What are our proposed guidelines’ primary advantages (if any) to support design adaptations in mHealth applications targeting chronic disease (Feel free to choose specific guideline numbers below if applicable)?
Are there any limitations or threats to the proposed guidelines (if they exist) for supporting design adaptations in mHealth applications targeting chronic disease (Feel free to choose specific guideline numbers below if applicable)?
What do you think could be done to accommodate these limitations/threats in the next version of the guidelines (Feel free to choose specific guideline numbers below if applicable)?
Please provide any other suggestions that you may have, e.g., about this research project, the developed guidelines etc. (Feel free to reference specific guideline numbers below if applicable.)